# Fair Foundation Models for Medical Image Analysis: Challenges and Perspectives

### v1

@@ 3.2 Data Curation
Although increasing data volume in parallel with neural network size can enhance model performance 92, 93, the relationship between data quantity and model improvement is not consistently linear. Recent research demonstrates that systematic data curation plays a critical role in both natural language processing and computer vision tasks 83, 58. Ensuring dataset fairness requires careful attention to multiple dimensions: diversity, global representativeness, equipment variability, disease representation, gender balance, and age distribution. Furthermore, data deduplication serves as an essential process for eliminating near-duplicate images, thereby reducing redundancy and enhancing dataset diversity and representativeness 58, 94.
@@
Sensitive attributes. The acquisition of metadata and labels in medical imaging presents substantial challenges, primarily due to the resource-intensive nature of manual curation given data volume and complexity. While fairness research frequently presumes the availability of demographic data, such information often remains inaccessible due to legal, ethical, and practical constraints 95. This limited access to demographic data creates significant obstacles for both data curation processes and the development of unbiased models 23. Current methodologies attempting to address fairness without demographic metadata face fundamental limitations, including systematic biases, accuracy limitations, technical barriers, and compromised transparency, thus emphasizing the critical need for rigorous evaluation protocols and explicit methodological guidelines 86.
@@
Scalable methods without sensitive attributes. FMs serve as highly effective feature extractors, enabling clustering-based approaches for automatic data curation that do not require sensitive attributes. This capability facilitates the enhancement of diversity and balance in large datasets 96, 23. When compared to uncurated data, features derived from these automatically curated datasets demonstrate superior performance 96. The clustering techniques developed through this approach serve a dual purpose: they not only improve data quality but also provide a systematic method for identifying potential biases in models 23. These findings establish foundation models as versatile tools that excel not only in feature extraction but also in crucial data management tasks, including bias assessment and automated curation.
@@
Data curation alone is insufficient. While data balancing contributes to model fairness, it does not fully eliminate inherent biases in AI systems. The integration of multiple data domains, particularly the combination of textual and visual data, can potentially introduce or intensify biases beyond those present in unimodal models 97, 98, 99. Consequently, achieving fair downstream behavior requires comprehensive mitigation strategies beyond data balancing alone 100. In the context of unimodal models, performance disparities in downstream tasks frequently correlate with class difficulty, where more challenging classes exhibit higher misclassification rates and decreased performance metrics 39.
@@ 4.1 Training
Fairness. Self-supervised learning models leveraging diverse, unlabeled datasets demonstrate enhanced fairness and inclusivity compared to supervised approaches, yielding outcomes characterized by increased robustness and reduced bias 93. However, FMs retain inherent biases despite these advantages 20, 22, 82. The mitigation of such biases necessitates targeted interventions within the training loop. This optimization process faces dual challenges: the substantial computational requirements and the inherent difficulty of acquiring large-scale datasets that exclude sensitive attributes.
@@
Bias Amplification in VLMs. VLMs face heightened challenges in bias mitigation due to their requirement for concurrent training on both visual and textual data modalities. Contemporary research demonstrates that multimodal training not only maintains existing societal biases 98, as evidenced in CLIP (Contrastive Language-Image Pretraining) architectures 97, but also intensifies these biases relative to single-modality systems 99.
@@ 4.1.1 Pre-training
Sensitive Attributes. The computational challenges inherent in vision and language models stem from power-law scaling relationships, wherein incremental performance improvements necessitate exponential increases in computational resources 92, 93. As established in the Data Creation section, the prevalent absence of metadata in large-scale datasets introduces significant complexities for training loop implementation in pre-trained FMs. These constraints fundamentally limit the application of fairness techniques during the training process. Furthermore, fairness methodologies developed for pre-trained FMs must demonstrate efficient scalability across expansive datasets and sophisticated architectural frameworks.
@@
Data selection. The incorporation of systematic data curation into the training loop can be achieved through active selection strategies, wherein computational priority is assigned to data elements that maximize task performance contributions 101. Research demonstrates that small curated models can facilitate the training of larger models through systematic identification of both straightforward and challenging image cases 102, 103. Implementation of these selection methodologies yields dual benefits: reduced training duration for pre-trained models and enhanced quality of resultant outputs.
@@
Loss. The Reducible Holdout Loss (RHO) employs a secondary model to identify three categories of data points: those that are learnable, those worth dedicating computational resources to learn, and those not yet acquired by the model 104. One established approach for enhancing model robustness and fairness involves the integration of sensitive attributes into loss functions, thereby promoting equitable outcomes across demographic groups 105. However, as examined in the Data Documentation section, the acquisition of sensitive attributes presents significant practical challenges. To address this limitation, alternative methodologies utilizing clustering-based data curation 96 offer potential proxy measures for sensitive attributes 23.
@@
Large Concept Models. Contemporary research in robust and unbiased model development prioritizes the prediction of representations within embedding spaces over traditional token-level forecasting approaches (e.g., word or pixel prediction) 106. Within this framework, a novel class of FMs, termed Large Concept Models (LCM), enables simultaneous training across multiple languages and modalities while maintaining scalability and minimizing bias 107. In the visual domain, the Image-based Joint-Embedding Predictive Architecture (I-JEPA) implements this embedding-centric approach, yielding substantial improvements in robustness, scalability, and computational efficiency relative to Masked Autoencoders (MAE) 52, 108.
@@ 4.1.2 Fine-tuning
Efficacy of Fine-tuning for Fairness Enhancement. Fine-tuning serves as a critical mechanism for aligning pre-trained FMs with targeted objectives. These models demonstrate exceptional capability in adapting to novel data distributions with minimal sample requirements 54. Fine-tuning effectively mitigates bias through increased model sensitivity to training distribution, particularly when employing balanced and curated datasets 100.
@@
Parameter-efficient Fine-tuning. Methodologies such as LoRa 109 and QLoRa 110 enhance the efficiency of low-resource fine-tuning through selective parameter modification, targeting only specific subsets within the base model architecture. The impact of these optimization techniques on bias mitigation remains an active area of investigation, with current research providing inconclusive evidence regarding their effects on model fairness 111, 25. Notably, alternative computational optimization strategies, including model pruning and differentially private training approaches, demonstrate increased bias manifestation within specific demographic subgroups 112, 113.
Mitigating Bias. Despite extensive research into bias mitigation strategies within deep learning frameworks, conventional fairness interventions demonstrate variable efficacy when applied to FMs 25. The limitations of these approaches extend beyond FMs applications, with traditional deep learning implementations often achieving only modest improvements in fairness metrics 114. Advanced data augmentation strategies, including AutoAug, Mixup, and CutMix methodologies, demonstrate promise in addressing challenging FMs scenarios 39. Furthermore, the integration of synthetically generated data during the fine-tuning process presents compelling evidence for enhanced fairness outcomes 90.
Blackbox FMs. The increasing commercialization of FMs and proliferation of their APIs has resulted in a distribution model where access is frequently restricted to embedding outputs, limiting direct interaction with model architectures. This constrained accessibility presents distinct challenges for bias mitigation in medical imaging applications. Recent research demonstrates, however, that effective bias elimination techniques can be implemented without requiring access to internal model parameters, offering viable solutions for both open-source and proprietary model frameworks 115.
@@ 4.2 Model Evaluation
Fairness Considerations and Metrics. Within deep learning research, fairness has become a critical consideration, with particular significance in medical imaging applications 28, 31, 5, 26, 22. The evaluation of fairness encompasses two fundamental methodological approaches: individual fairness, which requires consistent model outputs for similar inputs, and group fairness, which assesses model performance across demographic categories defined by sensitive attributes such as race and gender. Despite the widespread adoption of group fairness metrics in practical applications, a significant number of influential FMs studies in medical imaging have conducted evaluations without incorporating explicit fairness metrics 54, 53, 116, 117, thus leaving critical questions about potential biases and their clinical implications unexplored.
Utility-Fairness Trade-offs. The pursuit of bias minimization can inadvertently impact primary predictive performance, particularly affecting non-protected demographic groups, thereby introducing considerations of beneficence in model development. Establishing an optimal equilibrium between utility and fairness represents a fundamental challenge in FMs implementation 118. Contemporary research introduces quantitative methodologies for evaluating these trade-offs within specific prediction tasks and group fairness frameworks 119. Empirical evidence suggests that the optimization of utility-fairness balance depends predominantly on model adaptation strategies rather than initial model selection; furthermore, certain adaptation approaches demonstrate the potential to concurrently enhance both utility metrics and fairness measures 25.
@@
Robustness. A fundamental challenge in medical imaging applications lies in ensuring that fairness-aware models maintain their equitable performance when transitioning between different data distributions (A to B), particularly given the dynamic nature of population characteristics and deployment contexts 120. Although adversarial training techniques enhance overall model robustness, the improvements typically manifest unevenly across different classes 121, 122. Research indicates that fairness enhancement strategies can simultaneously strengthen model robustness 123. However, contemporary deep learning systems frequently demonstrate inadequate robustness to sensitive demographic attributes, specifically sex and gender variables 114. Within FMs frameworks, the direct deployment of models without fine-tuning procedures further compromises robustness to these demographic attributes, thus constraining their efficacy across diverse application scenarios 23.
@@
Data-Efficient Generalization. FMs demonstrate exceptional capabilities in data-efficient generalization, facilitating fine-tuning processes for medical imaging tasks with minimal labeled data requirements 54. This characteristic proves particularly advantageous for resource-constrained institutions implementing model deployment within their specific data environments. Nevertheless, substantial uncertainty persists regarding the implications of limited labeled data usage on fairness metrics in FMs applications. The development of ethical frameworks and bias mitigation strategies assumes critical importance in these contexts, especially given that institutions operating under resource constraints typically depend on restricted labeled data availability.
@@
Benchmark. Benchmarks are essential for assessing fairness in FMs development, providing teams with systematic evaluation tools. While fairness benchmarks exist for deep learning in medical imaging 114, 124, 125, 126, comprehensive benchmarks and libraries specifically designed for FMs remain lacking 25, 19. Developing such resources is crucial for holistic fairness evaluation, encompassing metrics, utility, utility-fairness trade-offs, robustness, and data-efficient generalization. Furthermore, FMs should undergo testing on unbiased data and pipelines, with evaluations conducted on distributions that reflect real-world applications to ensure valid and applicable fairness metrics 11.

### v2

@@ 3.2 Data Curation
Although increasing data volume in parallel with neural network size can enhance model performance [54, 55], the relationship between data quantity and model improvement is not consistently linear. Recent research demonstrates that systematic data curation plays a critical role in both natural language processing and computer vision tasks [39, 117]. Ensuring dataset fairness requires careful attention to multiple dimensions: diversity, global representativeness, equipment variability, disease representation, gender balance, and age distribution. Furthermore, data deduplication serves as an essential process for eliminating near-duplicate images, thereby reducing redundancy and enhancing dataset diversity and representativeness [117, 84]. In parallel, removing noisy samples, incomplete entries, and ambiguous abbreviations is crucial for mitigating hallucinations and achieve fair representations [61, 79].
@@
Protected attributes. The acquisition of metadata and labels in medical imaging presents substantial challenges, primarily due to the resource-intensive nature of manual curation given data volume and complexity. While fairness research frequently presumes the availability of demographic data, such information often remains inaccessible due to legal, ethical, and practical constraints [6]. This limited access to demographic data creates significant obstacles for both data curation processes and the development of unbiased models [123]. Current methodologies attempting to address fairness without demographic metadata face fundamental limitations, including systematic biases, accuracy limitations, technical barriers, and compromised transparency, thus emphasizing the critical need for rigorous evaluation protocols and explicit methodological guidelines [7].
Text and image. Recent research on VLMs highlights that data curation plays a more critical role in CLIP model performance than the architecture itself [160]. While enhancements such as the addition of sigmoid loss contribute to performance [166], the success of VLMs primarily depends on abundant high-quality data and substantial compute resources for scaling [141]. The process of curation leverages text metadata to construct more relevant datasets by selecting key terms, applying substring matching, and most importantly, balancing the dataset through sub-sampling based on chosen metadata [160].
@@
Scalable methods without protected attributes. FMs serve as highly effective feature extractors, enabling clustering-based approaches for automatic data curation that do not require protected attributes. This capability facilitates the enhancement of diversity and balance in large datasets [146, 123]. When compared to uncurated data, features derived from these automatically curated datasets demonstrate superior performance [146], for example by reducing a 4.44% gender imbalance to achieve an equal distribution of male and female samples in medical imaging [123]. The clustering techniques developed through this approach serve a dual purpose: they not only improve data quality but also provide a systematic method for identifying potential biases in models [123]. These findings establish FMs as versatile tools that excel not only in feature extraction but also in crucial data management tasks, including bias assessment and automated curation.
@@
Data curation alone is insufficient. While data balancing contributes to model fairness, it does not fully eliminate inherent biases in AI systems. The integration of multiple data domains, particularly the combination of textual and visual data, can potentially introduce or intensify biases beyond those present in unimodal models [59, 66, 19]. Consequently, achieving fair downstream behavior requires comprehensive mitigation strategies beyond data balancing alone [3]. In the context of unimodal models, performance disparities in downstream tasks frequently correlate with class difficulty, where more challenging classes exhibit higher misclassification rates and decreased performance metrics [36].
@@ 4.1 Training
Fairness. Self-supervised learning models leveraging diverse, unlabeled datasets demonstrate enhanced fairness and inclusivity compared to supervised approaches, yielding outcomes characterized by increased robustness and reduced bias [55]. However, FMs retain inherent biases despite these advantages [53, 144, 97]. For example, in a chest radiograph classification task, female patients experienced a 6.8–7.8% decrease in performance on the “no finding” label [52]. The mitigation of such biases necessitates targeted interventions within the training loop. This optimization process faces dual challenges: the substantial computational requirements and the inherent difficulty of acquiring large-scale datasets that exclude protected attributes.
@@
Bias. The lack of data representativeness, as illustrated in Figure 2, can cause models to perform disproportionately better on more prevalent or overrepresented data subsets, thereby limiting generalizability [36]. Models are known to internalize diverse biases related to complexity, shape, and other intrinsic features present in training data, which compromises their robustness [136, 82]. Hence, assembling datasets that comprehensively reflect the true population distribution is paramount to enhancing model robustness and mitigating bias-related performance degradation during training.
Bias Amplification in VLMs. Multimodal models not only perpetuate existing societal biases but also amplify them relative to single-modality systems [66, 59, 19]. The pairing of images and text reintroduces the well-documented problem of label bias [148, 26], and generating non-factual medical diagnoses and overconfidence in generated diagnoses [156]. Moreover, as shown in the Table 2, most multimodal datasets remain English-centric, which restricts their applicability in global contexts [12]. This limitation is reinforced by the fact that most multilingual solutions rely heavily on translation pipelines, wich could introduce semantic drift and cultural misalignment [40], in the medical domain this could lead to a misdiagnosis of local disease.
Synthetic data. Integrating synthetically generated data during training offers a promising approach to enhancing fairness in FMs. Evidence demonstrates that models trained exclusively on synthetic data can match or exceed the accuracy of models trained with real data [78, 111, 81]. For instance, in chest radiography applications, models trained with synthetic data achieved a 6.5% accuracy increase in downstream classification performance compared to models trained with real data [111]. Models that integrate medical findings with patient demographics prove essential for generating well-distributed synthetic datasets, as exemplified by RoentGen-v2 [111]. Model performance depends critically on synthetic data quality. Traditional high-quality generation methods require substantial computational resources. Emerging techniques such as flow matching and consistency models address this limitation through end-to-end architectures that reduce computational demands while maintaining generation quality [51, 135].
@@ 4.1.1 Pre-training
Protected attributes. The computational challenges inherent in vision and language models stem from power-law scaling relationships, wherein incremental performance improvements necessitate exponential increases in computational resources [54, 55]. As established in the Data Creation section, the prevalent absence of metadata in large-scale datasets introduces significant complexities for training loop implementation in pre-trained FMs. These constraints fundamentally limit the application of fairness techniques during the training process. Furthermore, fairness methodologies developed for pre-trained FMs must demonstrate efficient scalability across expansive datasets and sophisticated architectural frameworks.
@@
Data selection. The incorporation of systematic data curation into the training loop can be achieved through active selection strategies, wherein computational priority is assigned to data elements that maximize task performance contributions [130]. Research demonstrates that small curated models can facilitate the training of larger models through systematic identification of both straightforward and challenging image cases [48, 47]. Implementation of these selection methodologies yields dual benefits: reduced training duration for pre-trained models and enhanced quality of resultant outputs.
@@
Loss. The Reducible Holdout Loss (RHO) employs a secondary model to identify three categories of data points: those that are learnable, those worth dedicating computational resources to learn, and those not yet acquired by the model [107]. One established approach for enhancing model robustness and fairness involves the integration of protected attributes into loss functions, thereby promoting equitable outcomes across demographic groups [102]. However, as examined in the Data Documentation section, the acquisition of protected attributes presents significant practical challenges. To address this limitation, alternative methodologies utilizing clustering-based data curation [146] offer potential proxy measures for protected attributes [123].
@@
World Models. Contemporary research in robust and unbiased model development prioritizes the prediction of representations within embedding spaces over traditional token-level forecasting approaches (e.g., word or pixel prediction) [83]. Within this framework, a novel class of FMs, termed World Models enables simultaneous training across multiple languages and modalities while maintaining scalability and minimizing bias [138]. In the visual domain, the I-JEPA implements this embedding-centric approach, yielding substantial improvements in robustness, scalability, and computational efficiency relative to MAE [8, 90].
Vision Language World Models. Current VLMs, such as LLaVA-Med [86] and MedVInT [171], exhibit notable challenges including overconfidence in diagnostic outputs, privacy breaches, and the perpetuation of health disparities, as mention in Section 4.1 [156]. A promising approach to mitigate these issues involves integrating language and vision within world models equipped with a planning mechanism [27]. This mechanism operates by evaluating the consequences of an action, generating multiple potential outcomes, and predicting the optimal course that minimizes a cost function as assessed by a critic model [27, 83]. Crucially, fairness constraints can be embedded into this critic model. As this component is a self-supervised language model, establishing robust fairness principles in the language domain becomes fundamental to mitigating systemic bias throughout the entire vision-language architecture.
Knowledge Agglomeration. Knowledge agglomeration has recently emerged as a promising approach for pre-training. This approach leverages knowledge from existing FMs with diverse perspectives, such as CLIP [3], DINO [118], and SAM [80], to train improved models. Recent examples, like RADIO [62], demonstrate that models trained via knowledge agglomeration can surpass their teacher’s performance. In the medical imaging domain, models such as MedSAM [98] and MedCLIP [149] can be combined to enhance performance. Crucially, this approach allows integrating models with fairness-aware representations alongside those optimized for general performance, jointly improving subgroup representation while balancing the fairness-utility trade-off (see Section 4.2 for a detailed discussion of utility-fairness trade-offs).
@@ 4.1.2 Fine-tuning
Efficacy of Fine-tuning for Fairness Enhancement. Fine-tuning serves as a critical mechanism for aligning pre-trained FMs with targeted objectives. These models demonstrate exceptional capability in adapting to novel data distributions with minimal sample requirements [9]. Fine-tuning effectively mitigates bias through increased model sensitivity to training distribution, particularly when employing balanced and curated datasets [3].
@@
Parameter-efficient Fine-tuning. Methodologies such as LoRa [64] and QLoRa [41] enhance the efficiency of low-resource fine-tuning through selective parameter modification, targeting only specific subsets within the base model architecture. The impact of these optimization techniques on bias mitigation remains an active area of investigation, with current research providing inconclusive evidence regarding their effects on model fairness [42, 71]. Notably, alternative computational optimization strategies, including model pruning and differentially private training approaches, demonstrate increased bias manifestation within specific demographic subgroups [140, 10].
Mitigating Bias. Despite extensive research into bias mitigation strategies within deep learning frameworks, conventional fairness interventions demonstrate variable efficacy when applied to FMs [71]. The limitations of these approaches extend beyond FMs applications, with traditional deep learning implementations often achieving only modest improvements in fairness metrics [176]. Advanced data augmentation strategies, including AutoAug, Mixup, and CutMix methodologies, demonstrate promise in addressing challenging FMs scenarios [36]. Furthermore, the integration of synthetically generated data during the fine-tuning process presents compelling evidence for enhanced fairness outcomes [81].
Blackbox FMs. The increasing commercialization of FMs and proliferation of their APIs has resulted in a distribution model where access is frequently restricted to embedding outputs, limiting direct interaction with model architectures. This constrained accessibility presents distinct challenges for bias and hallucinations mitigations in medical imaging applications. Recent research demonstrates, however, that effective bias and hallucinations eliminations techniques can be implemented without requiring access to internal model parameters, offering viable solutions for both open-source and proprietary model frameworks [70, 101].
@@ 4.2 Model Evaluation
Fairness Considerations and Metrics. Within deep learning research, fairness has become a critical consideration, with particular significance in medical imaging applications [30, 161, 129, 133, 144]. The evaluation of fairness encompasses two fundamental methodological approaches: individual fairness, which requires consistent model outputs for similar inputs, and group fairness, which assesses model performance across demographic categories defined by protected attributes such as race and gender. Despite the widespread adoption of group fairness metrics in practical applications, a significant number of influential FMs studies in medical imaging have conducted evaluations without incorporating explicit fairness metrics [9, 174, 96, 173], thus leaving critical questions about potential biases and their clinical implications unexplored.
Utility-Fairness Trade-offs. In real-world scenarios, models may use protected attributes as shortcuts for specific tasks, introducing bias [162]. However, bias does not always stem from such shortcuts; other factors like intensity-based and morphology-based effects can also influence model behavior [136]. This bias complexity creates a utility-fairness trade-off, where achieving statistical parity can reduce utility for some groups [172, 150]. Identifying the optimal trade-off point is essential to balance both metrics effectively [38, 119].
Hallucinations. Medical hallucinations arise in specialized tasks such as diagnostic reasoning, employing domain-specific terminology that masks critical inaccuracies[79]. Expert scrutiny remains essential for detecting these plausible yet incorrect outputs, which can delay appropriate interventions and compromise patient safety [79]. Large language model taxonomies typically categorize hallucinations into factual errors, outdated references, spurious correlations, incomplete reasoning chains, and fabricated sources across text and multimodal contexts [79]. Detecting and quantifying hallucinations is critical for ensuring model safety and reliability [29, 79, 57]. Exploring the relationship between protected attributes and hallucinations represents an emerging research area that advances fairness evaluations and addresses bias in generative models [111, 79]. In this context, rigorous quality checks of generated images are essential to remove hallucinations and mitigate bias; for example, in a dataset of 623,712 prompts, 58,558 (9.4%) failed quality control, predominantly due to violations of race-related criteria [111].
@@
Robustness. A fundamental challenge in medical imaging applications lies in ensuring that fairness-aware models maintain their equitable performance when transitioning between different data distributions (A to B), particularly given the dynamic nature of population characteristics and deployment contexts [131]. Although adversarial training techniques enhance overall model robustness, the improvements typically manifest unevenly across different classes [159, 99]. Research indicates that fairness enhancement strategies can simultaneously strengthen model robustness [112]. However, contemporary deep learning systems frequently demonstrate inadequate robustness to protected attributes, specifically sex and gender variables [176]. Within FMs frameworks, the direct deployment of models without fine-tuning procedures further compromises robustness to these demographic attributes, thus constraining their efficacy across diverse application scenarios [123].
@@
Data-Efficient Generalization. FMs demonstrate exceptional capabilities in data-efficient generalization, facilitating fine-tuning processes for medical imaging tasks with minimal labeled data requirements [9]. This characteristic proves particularly advantageous for resource-constrained institutions implementing model deployment within their specific data environments. Nevertheless, substantial uncertainty persists regarding the implications of limited labeled data usage on fairness metrics in FMs applications. The development of ethical frameworks and bias mitigation strategies assumes critical importance in these contexts, especially given that institutions operating under resource constraints typically depend on restricted labeled data availability.
@@
Benchmark. Benchmarks are essential for assessing fairness in FMs development, providing teams with systematic evaluation tools. While fairness benchmarks exist for deep learning in medical imaging [176, 167, 175, 45], comprehensive benchmarks and libraries specifically designed for FMs remain lacking [71, 77]. Developing such resources is crucial for holistic fairness evaluation, encompassing metrics, utility, utility-fairness trade-offs, robustness, and data-efficient generalization. Furthermore, FMs should undergo testing on unbiased data and pipelines, with evaluations conducted on distributions that reflect real-world applications to ensure valid and applicable fairness metrics [94]. Benchmarks such as Med-HallMark, which introduces detection methods including MediHallDetector, provide systematic frameworks for evaluating and mitigating hallucinations in medical applications [29].
The Fairness Benchmarking for Medical Imaging Foundation Models (FairMedFM) [71] provides a comprehensive framework for comparing models, techniques, and datasets, covering most of the aspects discussed here. Given the importance of fairness and the lack of protected attributes in many datasets (see Table 2), we encourage both new and existing datasets to collect this information. Even if such metadata is limited to the test set or a reduced patient cohort, it should accurately reflect the dataset distribution to support fairness evaluation and model development, as demonstrated by datasets like Harvard-FairVLMed [97]. Legacy datasets such as PMC-15M [170], which are significant for VLMs research, could be enhanced by extracting demographic metadata such as age and sex from captions. Methods for dataset curation like MetaCLIP [160] can facilitate this process, improving both evaluation and model training. In cases where protected attributes cannot be obtained, we recommend techniques that create groups approximating demographic attributes to facilitate fairness analysis [123, 146].